Find the three consecutive even numbers such that sum of the first and the last numbers exceeds the second number by 10



Answer by mananth(16949)

Find the three consecutive even numbers such that sum of the first and the last numbers exceeds the second number by 10
..
let the numbers be x,x+2,x+4
..
x+x+4 = x+2+10
2x+4 = x+12
2x-x 12-4
x=8
..
the numbers are 8,10,12
CHECK
8+12= 10+10